OpenStudy (anonymous):

Slope of the tangent line to the polar curve r=sin(6theta). Theta =pi/12. Find slope

OpenStudy (anonymous):

Please help

OpenStudy (irishboy123):

you are looking for the slope in x-y coordinates you need this: \(\huge \frac{dy}{dx} = \frac{\frac{dr}{d \theta} sin \ \theta + r \ cos \ \theta} {\frac{dr}{d \theta} cos \theta - r sin \ \theta}\) \(\large \frac{dr}{d \theta} = \frac{d}{d \theta} (sin6\theta) = 6 cos 6 \theta\) \(\large \theta = \frac{\pi}{12} \implies \frac{dr}{d \theta} = 6 cos (6 *(\frac{\pi}{12})) =6 cos (\pi/2 ) = 0\) \(\large r(\frac{\pi}{12}) = sin(6 \theta) = sin(6*\frac{\pi}{12}) = sin( \pi / 2) = 1\) \(\large \frac{dy}{dx} = -cot (\frac{\pi}{12}) \)

OpenStudy (anonymous):

So what is the slope?

OpenStudy (australopithecus):

You need the formula \[\frac{dy}{dx} = \frac{\frac{dr}{d \theta} \sin(\theta) + r*\cos(\theta)}{\frac{dr}{d \theta}\cos(\theta) - r*\sin(\theta)}\] so take the derivative of sin(6theta) and plug into the formula for example if I had 4 The derivative would be 0 so I would plug that into my formula: (0*sin(theta) + 4cos(theta) )/(0*cos(theta) -4sin(theta)) Simplify then plug Theta =pi/12 into your formula and solve and you should get your derivative

OpenStudy (australopithecus):

in my example r = 4

OpenStudy (australopithecus):

This video explains the method nicely https://www.youtube.com/watch?v=GkhOx4hUssA

OpenStudy (anonymous):

What would it be for me

OpenStudy (irishboy123):

-3.73 i make it steep !

OpenStudy (australopithecus):

I will give you an example one second

OpenStudy (anonymous):

Thank u

OpenStudy (irishboy123):

again \( \theta\) can be blagged because \(\theta = \pi / 12\) sits on the extreme of a petal and on the circle [of radius 1]

OpenStudy (australopithecus):

Here is the explicit answer just hit approximate for result and you will see the result I recommend you check it out

OpenStudy (mathmate):

\(r=sin(6\theta)~ and~ \theta=\pi/12, so~ r=sin(6\pi/12)=sin(\pi/2)=1\)